# Weak Cosmic Censorship and Weak Gravity Conjectures in CFT Thermodynamics

### v1

@@ Abstract
In this paper, we explore the intriguing interplay between fundamental theoretical physics concepts within the context of charged black holes. Specifically, we focus on the consistency of the weak gravity conjecture (WGC) and weak cosmic censorship conjecture (WCCC) in the thermodynamics of conformal field theory (CFT), and restricted phase space (RPS) for AdS Reissner-Nordström black holes with a perfect fluid dark matter (RN-PFDM). The WGC ensures that gravity remains the weakest force in the system. Meanwhile, the WCCC addresses the cosmic censorship problem by preventing the violation of fundamental physical laws near the black hole singularity. First, we analyze the RN black hole’s free energy in both spaces, revealing a distinctive swallowtail pattern indicative of a first-order phase transition when certain free parameter conditions are met. We explore the WGC across different phase spaces, emphasizing the need for certain conditions in EPST, RPST, and CFT. We demonstrate that PFDM $\gamma$ and the radius of AdS $l$ have a vital role in proving the satisfaction of the WGC. Also, these values have a linear relation with the range compatibility of WGC. The range of compatibility for WGC in RPST and EPST is the same while for CFT, this range is larger than EPST, and RPST. It means somehow the WGC and CFT are more consistent. The WCCC was examined at the critical juncture, confirming its validity with $f(M_{c},Q_{c},r_{c})=0$. We conclude that the WGC is supported at the critical point of black holes, and the WCCC is also maintained, demonstrating the robustness of these conjectures within the critical ranges of black hole parameters.
@@ 1 Introduction
The exploration of black hole thermodynamics, initially spurred by the work of Hawking and Bekenstein, delves into the thermal characteristics of these cosmic entities from a thermodynamic physics perspective[1, 2, 3]. This field posits a correlation between surface gravity and the event horizon’s area with temperature and entropy, respectively, while often treating a black hole’s mass as its internal energy. Despite the progress achieved, questions remain, particularly regarding the microscopic contributors to black hole entropy. The principles of holography and the AdS/CFT duality are instrumental in examining various black hole properties, where a state of an AdS black hole in the bulk corresponds to a thermal state in the dual field theory[4, 5]. Additionally, the thermodynamics of black holes can be analyzed through the lens of different symmetries within a gravitational model[6]. Hawking and Page’s advanced research on AdS black holes revealed a phase transition between such black holes accompanied by radiation and their thermal AdS counterparts[7]. The thermodynamic properties of black holes, especially within the framework of extended phase space thermodynamics, have been extensively discussed, where the negative cosmological constant is equated with pressure[8, 9, 10, 11, 12]. This perspective has led to the inclusion of other model parameters as novel thermodynamic variables[13, 14].
@@
Furthermore, the use of P-V variables suggests that AdS black holes could function as heat engines[15, 16]. A recent theoretical development has introduced central charge and chemical potential as a new pair of conjugate thermodynamic variables, enhancing our understanding of black hole thermodynamics. In the corresponding dual theory, the central charge is linked to the square of the number of colors[17, 18, 19]. In the realm of AdS/CFT correspondence, a novel thermodynamic framework known as restricted phase space thermodynamics (RPST) has emerged, as introduced by Gao et al. This paradigm redefines black hole masses as a form of internal energy. Unlike the extended phase space thermodynamics (EPST) that utilizes pressure and volume, RPST employs a distinct set of conjugate thermodynamic variables: the chemical potential, denoted as µ, and the central charge, C. These variables represent color susceptibility and the degrees of freedom within the conformal field theory (CFT), respectively[18, 19, 20]. The central charge, C, plays a role analogous to the particle count in the statistical physics of conventional matter. It is mathematically expressed in relation to Newton’s constant, G, and the anti-de Sitter (AdS) curvature radius, through the formula $C=\frac{l^{d-2}}{G}$ Here, $d$ signifies the dimensionality of the AdS spacetime. The RPST framework has been applied to explore the homogeneity of the Smarr relation, various thermodynamic processes, and phase transitions across different black hole configurations such as Reissner-Nordström-AdS (RN-AdS), Kerr-AdS, Bañados-Teitelboim-Zanelli (BTZ), and Taub-NUT within the AdS context[20, 21, 22, 23, 24, 25].
@@
The holographic principle has revolutionized our understanding of space and information, suggesting that the information contained within a volume of space can be represented on a lower-dimensional boundary. This concept has given rise to the field of holographic black hole chemistry. It posits a duality between gravitational dynamics in higher-dimensional Anti-de Sitter (AdS) space and quantum field theory on its boundary, a relationship known as the AdS/CFT correspondence. Through this lens, we can explore black hole thermodynamics by treating the cosmological constant as a variable akin to thermodynamic pressure in AdS space or a scaling factor on the boundary. This innovative approach has led to the development of black hole chemistry, which can define thermodynamic quantities like enthalpy, Gibbs free energy, and chemical potential for black holes[26, 27, 28, 29]. It allows probing of the properties and phase transitions of black holes in AdS space through holographic duality. Notably, it provides insights into phenomena such as the Hawking-Page transition, which reflects a phase shift from thermal AdS space to a large AdS black hole, analogous to the confinement/deconfinement transition in conformal field theory. One of the intriguing outcomes is that AdS black holes adhere to a generalized Smarr relation, linking their mass, area, charge, angular momentum, and pressure. However, integrating the first law of black hole thermodynamics with the thermodynamics of its holographic dual presents challenges. This research is particularly focused on how the variation of the cosmological constant ($\Lambda$) correlates with changes in the central charge (C) and the volume (V) of the boundary CFT. This relationship necessitates a revision of the thermodynamic laws to account for the volume and its associated thermodynamic pressure[26, 27, 28, 29]. The possibility of varying fundamental constants, like Newton’s constant (G) and the cosmological constant, hints at more profound underlying theories. researchers addressed this by demonstrating that altering G alongside $\Lambda$ maintains the constancy of the central charge C in the dual CFT. Consequently, the first law can be reformulated to include a new thermodynamic volume and chemical potential, bridging holographic and bulk thermodynamics through the central charge duality. Their work also uncovered phase structures with a critical central charge value, beyond which the phase diagram shows a swallowtail pattern indicative of a first-order phase transition. By treating G as a variable, we can incorporate the central charge into the first law and extend the thermodynamic discussion to various black holes. In essence, black hole chemistry reveals that AdS black holes experience phase transitions mirroring those found in conventional thermodynamic systems. Yet, the extended thermodynamics’ holographic interpretation remained elusive for some time. Initial theories proposed that the volume-pressure term in AdS should correspond to a chemical potential-central charge term in the dual CFT, where the central charge is C and the chemical potential is its thermodynamic conjugate[26, 27, 28, 29].
The Swampland Program is an initiative in theoretical physics that seeks to identify effective low-energy theories which are not compatible with quantum gravity. This contrasts with the ”string theory landscape,” which includes theories known to be compatible with string theory, a consistent quantum theory of gravity. The Swampland encompasses theories that appear consistent but lack a consistent ultraviolet completion when gravity is added. The program was initiated by Cumrun Vafa and aims to delineate the boundaries of the quantum gravity landscape by identifying universal principles shared among all theories compatible with gravitational UV completion[30, 31, 32, 33, 34, 35, 36, 37, 38]. The WGC posits that in any theory of quantum gravity, gauge forces must mediate interactions that are stronger than gravity for some particles. This conjecture has deep connections across various areas of physics and mathematics, suggesting that gravity should be the weakest force in any consistent theory of quantum gravity. It has been reviewed extensively and is supported by evidence from string theory examples. The Swampland Program or the WGC is significant in the ongoing effort to understand the fundamental principles that govern our universe at the quantum level. They have implications for particle physics, cosmology, general relativity, and mathematics, guiding researchers in exploring the limits and possibilities of theoretical physics. To study an example of applications of the swampland program in inflation, black holes, thermodynamics, and black brane, you can see in[39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102].
In this exploration, we examine the nuanced interplay between the WGC and the WCCC in the realm of charged black holes. Our focus is on the (RN) black holes that are encased within PFDM, considering diverse spatial frameworks like RPS and CFT. This analysis aims to deepen our understanding of the theoretical underpinnings that govern these cosmic entities. The WCCC, introduced by Roger Penrose, is a pivotal concept in theoretical physics that addresses the potential issues arising from singularities within general relativity. The WCCC posits that singularities resulting from gravitational collapse are invariably concealed behind event horizons, thereby maintaining the causal structure and predictability of spacetime. This conjecture ensures that the ”naked” singularities, which could expose the inner workings of black holes, do not disrupt the observable universe. In contrast, the WGC suggests that in any consistent theory of quantum gravity, there must exist particles for which the electromagnetic force is stronger than their gravitational pull. This implies that black holes should be able to decay into these particles, preventing the formation of stable remnants and ensuring the discharge of black holes. The Reissner–Nordström (RN) black hole, a solution to the Einstein–Maxwell equations for a charged black hole, presents a unique challenge to these conjectures. If the charge $Q$ of the black hole exceeds its mass $M$ ($Q>M$), the RN solution would contravene the WCCC by revealing a naked singularity to distant observers. Conversely, when the RN black hole approaches an extremal state ($Q=M$), it suggests the existence of decay products with a charge-to-mass ratio exceeding unity. While this satisfies the WGC, it simultaneously breaches the WCCC, as these decay products cannot form black holes but are instead predicted to be elementary particles. This apparent contradiction between the WGC and the WCCC in the context of RN black holes is one of the intriguing puzzles at the intersection of quantum theory and gravitational physics[103, 104, 105].
The WGC and the WCCC are two distinct ideas in theoretical physics, each addressing different aspects of fundamental physics. While both conjectures involve the term ”weak” and pertain to fundamental aspects of gravity and spacetime, they are distinct in their focus and implications. The WGC primarily concerns the relative strengths of forces in a quantum gravity framework, while the WCCC deals with the behavior of singularities and event horizons in gravitational collapse scenarios within classical general relativity. Combining these concepts, one might speculate about potential connections between the WGC and cosmic censorship, such as whether the WGC could have implications for cosmic censorship scenarios. Building on the concepts previously discussed, we delve into the intricate relationship between the WGC and the WCCC within the framework of charged black holes, specifically the RN AdS black hole surrounded by PFDM. We examine this relationship from a thermodynamic standpoint, considering various theoretical spaces such as RPS and CFT by mentioning an important point that means. The Mild WGC posits that gravity is the weakest force, yet it accommodates a wider spectrum of theoretical possibilities, making it less stringent than its counterpart, the Strong WGC. The latter is more narrowly focused on particles, asserting not just that gravity must be weaker but also stipulating the existence of particles that strictly adhere to this inequality. When considering black holes, the WCCC maintains that singularities should remain veiled behind event horizons, safeguarding the causal structure of spacetime. The interplay between the WGC and WCCC is subtle and complex. Our investigations have revealed that under specific conditions, such as the incorporation of PFDM, these conjectures can indeed coexist harmoniously. This concord is vital as it ensures that our theoretical models remain consistent and do not yield contradictory predictions about the thermodynamic behavior of black holes. By demonstrating compatibility under these conditions, we contribute to a more coherent understanding of the fundamental principles governing black hole physics. The paper is organized as follows.
In Section 2, we conduct a comprehensive study of the RN AdS black hole in the presence of PFDM. Sections 3 and 4 are dedicated to examining the thermodynamics of the RN AdS black hole surrounded by PFDM within the frameworks of RPST and CFT, respectively. We also analyze the thermodynamic parameters in these two spaces. Section 5 explores the compatibility of the WGC and the WCCC in both spaces, delving into the areas where these conjectures align. Finally, in Section 6, we briefly summarize our findings.
2 Weak gravity conjecture
@@
The Weak Gravity Conjecture is a theoretical framework in the field of quantum gravity that suggests gravity should be the weakest force in any consistent theory of quantum gravity. It posits that for any gauge force, there must exist some particles for which the gauge force mediates interactions stronger than gravity. This conjecture has profound implications across various domains of physics and mathematics, and it has been a subject of significant interest and debate within the theoretical physics community. The WGC is not without its criticisms and potential weak points. Some of the criticisms include that the conjecture is based on qualitative arguments rather than a rigorous mathematical proof [51]. There are several variations of the conjecture, and it’s not clear which, if any, is the correct one. The conjecture has been challenged by certain string theory examples and other theoretical models that do not seem to satisfy its conditions. The WGC has made several predictions, including: The existence of particles with specific mass-to-charge ratios ensures gravity remains the weakest force. Constraints on the properties of black holes, particularly regarding their charge and mass. Implications for the stability of cosmic structures and the behavior of fundamental particles.
@@
The WGC has interesting connections with thermodynamics, particularly in the context of black hole physics. It has been suggested that the conjecture could be derived from thermodynamic principles applied to black holes, such as entropy considerations and the behavior of quasinormal modes [72, 75]. The implications of the WGC can extend to various spacetime geometries, including extended, restricted, and CFT spaces: The conjecture has been explored in the context of theories with extra dimensions, such as string theory, where it places constraints on the geometry of the compactified dimensions. In more constrained spacetime geometries, like those with a limited number of dimensions or symmetries, the WGC still implies the existence of particles that satisfy its mass-to-charge ratio condition. In the AdS/CFT correspondence, the WGC translates into inequalities involving the dimension and charge of operators, as well as central charges in the dual CFTs. However, these translations are still speculative and may not apply universally across all CFTs [52, 53]. In summary, the WGC is a conjecture that aims to provide insights into the nature of gravity in quantum gravity theories. It has sparked a wide range of research, leading to discussions about its validity, implications, and connections to other physical principles. While it has faced criticism and challenges, it continues to be a pivotal concept in theoretical physics, driving exploration into the fundamental forces of nature. The WGC has far-reaching implications that extend beyond the realm of theoretical physics into cosmology, black hole thermodynamics, and the broader landscape of string theory. In cosmology, the WGC can influence our understanding of the early universe, particularly during the inflationary period. The conjecture suggests that inflationary models involving scalar fields should be coupled to gauge fields in a way that satisfies the WGC criteria [54]. This can lead to predictions about the spectrum of primordial fluctuations, potentially observable in the Cosmic Microwave Background (CMB) data. String theory provides a fertile ground for testing the WGC. The conjecture has been used to argue for the existence of certain particles in the string spectrum that have implications for the stability of extra dimensions and the nature of supersymmetry breaking. A proof of the WGC within the context of perturbative string theory has been proposed, which relates the black hole extremality bound to long-range forces calculated on the worldsheet [55]. The thermodynamics of black holes offers a unique testing ground for the WGC. The conjecture has been studied in the context of 4D Einstein-Maxwell-dilaton theory, where it places constraints on the Wilson coefficients of the theory, which are satisfied under certain assumption about the UV completion. This study bridges the gap between quantum field theory and gravitational phenomena, providing insights into the entropy and stability of black holes [56]. The WGC may serve as a bridge between quantum mechanics and cosmic-scale physics. By examining the thermodynamics of black holes and their connection with the WGC, we can explore how quantum-level interactions manifest at cosmic scales. This could lead to a better understanding of the holographic principle and the nature of spacetime itself. The phenomenological implications of the WGC are vast. It can affect predictions for particle accelerators, inform the search for dark matter candidates, and influence models of dark energy. The conjecture’s requirement for the existence of certain particles could lead to observable signatures in high-energy physics experiments. In summary, the WGC is not just a theoretical curiosity but a principle that can potentially offer solutions to existing problems in cosmology and string theory, provide a framework for understanding black hole thermodynamics, and open new avenues for research in the phenomenology of quantum gravity theories. The ongoing dialogue between theory and observation will continue to refine our understanding of the WGC and its implications across the cosmos. Considering all the mentioned concepts, we must now explore other applications of the WGC within various cosmological frameworks. This exploration aims to address potential issues and provide evidence supporting both the conjecture itself and string theory. By examining black hole thermodynamics and its relationship with the WGC, we can establish a link from quantum to cosmic scales. This not only allows for a review of existing concepts but also enables the results obtained to unveil new perspectives on the implications and phenomenology associated with the WGC.

### v2

@@ Abstract
In this paper, we explore the intriguing interplay between fundamental theoretical physics concepts within the context of charged black holes. Specifically, we focus on the consistency of the weak gravity conjecture (WGC) and weak cosmic censorship conjecture (WCCC) in the thermodynamics of conformal field theory (CFT), and restricted phase space thermodynamics (RPST) for AdS Reissner-Nordström black holes with a perfect fluid dark matter (RN-PFDM). The WGC ensures that gravity remains the weakest force in the system. Meanwhile, the WCCC addresses the cosmic censorship problem by preventing the violation of fundamental physical laws near the black hole singularity. First, we analyze the RN black hole’s free energy in both spaces, revealing a distinctive swallowtail pattern indicative of a first-order phase transition when certain free parameter conditions are met. We explore the WGC across different phase spaces, emphasizing the need for certain conditions in extended phase space thermodynamics (EPST), RPST, and CFT. We demonstrate that PFDM parameter $\gamma$ and the radius of AdS $l$ have a vital role in proving the satisfaction of the WGC. Also, these values have a linear relation with the range compatibility of WGC. The range of compatibility for WGC in RPST and EPST is the same while for CFT, this range is larger than EPST, and RPST. It means somehow the WGC and CFT are more consistent. The WCCC was examined at the critical juncture, confirming its validity in critical points. We conclude that the WGC is supported at the critical point of black holes, and the WCCC is also maintained, demonstrating the robustness of these conjectures within the critical ranges of black hole parameters.
@@ 1 Introduction
The principles of holography and the AdS/CFT duality are instrumental in examining various black hole properties, where a state of an AdS black hole in the bulk corresponds to a thermal state in the dual field theory[1, 2, 3, 4, 5]. The thermodynamics of black holes can be analyzed through the lens of different symmetries within a gravitational model[6]. Hawking and Page’s advanced research on AdS black holes revealed a phase transition between such black holes accompanied by radiation and their thermal AdS counterparts[7]. The thermodynamic properties of black holes, especially within the framework of extended phase space thermodynamics, have been extensively discussed, where the negative cosmological constant is equated with pressure[8, 9, 10, 11, 12]. This perspective has led to the inclusion of other model parameters as novel thermodynamic variables[13, 14, 15, 16].
@@
A recent theoretical development has introduced central charge and chemical potential as a new pair of conjugate thermodynamic variables, enhancing our understanding of black hole thermodynamics. In the corresponding dual theory, the central charge is linked to the square of the number of colors[17, 18, 19]. In the realm of AdS/CFT correspondence, a novel thermodynamic framework known as RPS has emerged, as introduced by Gao et al[21, 22]. This paradigm redefines black hole masses as a form of internal energy. Unlike the EPST which utilizes pressure and volume, RPST employs a distinct set of conjugate thermodynamic variables: the chemical potential, denoted as $\mu$, and the central charge, C. These variables represent color susceptibility and the degrees of freedom within the conformal field theory (CFT), respectively[18, 19, 20]. The central charge, C, plays a role analogous to the particle count in the statistical physics of conventional matter. It is mathematically expressed in relation to Newton’s constant, G, and the anti-de Sitter (AdS) curvature radius, through the formula $C=\frac{l^{d-2}}{G}$. Here, $d$ signifies the dimensionality of the AdS spacetime. With respect to demonstrating that altering G alongside $\Lambda$ maintains the constancy of the central charge C in the dual CFT, so can consequently, the first law can be reformulated to include a new thermodynamic volume and chemical potential, bridging holographic and bulk thermodynamics through the central charge duality. The RPST framework has been applied to explore the homogeneity of the Smarr relation, various thermodynamic processes, and phase transitions across different black holes[20, 21, 23, 24, 25, 26].
@@
The properties and phase transitions of black holes in AdS space through holographic duality provide insights into phenomena such as the Hawking-Page transition, which reflects a phase shift from thermal AdS space to a large AdS black hole, analogous to the confinement/deconfinement transition in conformal field theory. One of the intriguing outcomes is that AdS black holes adhere to a generalized Smarr relation, linking their mass, area, charge, angular momentum, and pressure. However, integrating the first law of black hole thermodynamics with the thermodynamics of its holographic dual presents challenges. Researchers are particularly focused on how the variation of the cosmological constant ($\Lambda$) correlates with changes in the central charge (C) and the volume (V) of the boundary CFT[27, 28, 29, 30]. They also uncovered phase structures with a critical central charge value, beyond which the phase diagram shows a swallowtail pattern indicative of a first-order phase transition. By treating G as a variable, we can incorporate the central charge into the first law and extend the thermodynamic discussion to various black holes. Yet, the extended thermodynamics’ holographic interpretation remained elusive for some time. Initial theories proposed that the volume-pressure term in AdS should correspond to a chemical potential-central charge term in the dual CFT, where the central charge is C and the chemical potential is its thermodynamic conjugate[27, 28, 29, 30].
@@
The swampland program is an initiative in theoretical physics that seeks to identify effective low-energy theories that are compatible with quantum gravity. The swampland encompasses theories that appear consistent but lack a consistent ultraviolet completion when gravity is added[31, 32, 33, 34, 35, 36, 37, 38, 39]. The Weak Gravity Conjecture is a theoretical framework in the field of quantum gravity that suggests gravity should be the weakest force in any consistent theory of quantum gravity. It posits that for any gauge force, there must exist some particles for which the gauge force mediates interactions stronger than gravity[31, 32, 33, 34, 35]. The WGC is not without its criticisms and potential weak points. Some of the criticisms include that the conjecture is based on qualitative arguments rather than a rigorous mathematical proof [40]. The WGC has made several predictions, including: The existence of particles with specific mass-to-charge ratios ensures gravity remains the weakest force[31, 32, 33, 34, 35]. Constraints on the properties of black holes, particularly regarding their charge and mass. Implications for the stability of cosmic structures and the behavior of fundamental particles[31, 32, 33, 34, 35].
@@
The WGC has interesting connections with thermodynamics, particularly in the context of black hole physics. It has been suggested that the conjecture could be derived from thermodynamic principles applied to black holes, such as entropy considerations and the behavior of quasinormal modes [41, 42]. The implications of the WGC can extend to various spacetime geometries, including extended, restricted, and CFT spaces: The conjecture has been explored in the context of theories with extra dimensions, such as string theory, where it places constraints on the geometry of the compactified dimensions[31, 32, 33, 34, 35]. In more constrained spacetime geometries, like those with a limited number of dimensions or symmetries, the WGC still implies the existence of particles that satisfy its mass-to-charge ratio condition[31, 32, 33, 34, 35]. In the AdS/CFT correspondence, the WGC translates into inequalities involving the dimension and charge of operators, as well as central charges in the dual CFTs. However, these translations are still speculative and may not apply universally across all CFTs [43, 44]. The WGC has far-reaching implications that extend beyond the realm of theoretical physics into cosmology, black hole thermodynamics, and the broader landscape of string theory. In cosmology, the WGC can influence our understanding of the early universe, particularly during the inflationary period. The conjecture suggests that inflationary models involving scalar fields should be coupled to gauge fields in a way that satisfies the WGC criteria [45]. This can lead to predictions about the spectrum of primordial fluctuations, potentially observable in the Cosmic Microwave Background (CMB) data. String theory provides a fertile ground for testing the WGC. The conjecture has been used to argue for the existence of certain particles in the string spectrum that have implications for the stability of extra dimensions and the nature of supersymmetry breaking. A proof of the WGC within the context of perturbative string theory has been proposed, which relates the black hole extremality bound to long-range forces calculated on the worldsheet [46]. The thermodynamics of black holes offers a unique testing ground for the WGC. The conjecture has been studied in the context of 4D Einstein-Maxwell-dilaton theory, where it places constraints on the Wilson coefficients of the theory, which are satisfied under certain assumptions about UV completion. This study bridges the gap between quantum field theory and gravitational phenomena, providing insights into the entropy and stability of black holes [47]. The WGC may serve as a bridge between quantum mechanics and cosmic-scale physics. By examining the thermodynamics of black holes and their connection with the WGC, we can explore how quantum-level interactions manifest at cosmic scales. This could lead to a better understanding of the holographic principle and the nature of spacetime itself. The phenomenological implications of the WGC are vast. It can affect predictions for particle accelerators, inform the search for dark matter candidates, and influence models of dark energy. The conjecture’s requirement for the existence of certain particles could lead to observable signatures in high-energy physics experiments. For more study, you can see in[48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106].
In this exploration, we examine the nuanced interplay between the WGC and the WCCC in the realm of charged black holes. Our focus is on the (RN) black holes that are encased within PFDM, considering diverse spatial frameworks like RPS and CFT. This analysis aims to deepen our understanding of the theoretical underpinnings that govern these cosmic entities. The WCCC posits that singularities resulting from gravitational collapse are invariably concealed behind event horizons, thereby maintaining the causal structure and predictability of spacetime. This conjecture ensures that the ”naked” singularities, which could expose the inner workings of black holes, do not disrupt the observable universe. In contrast, the WGC suggests that in any consistent theory of quantum gravity, there must exist particles for which the electromagnetic force is stronger than their gravitational pull. This implies that black holes should be able to decay into these particles, preventing the formation of stable remnants and ensuring the discharge of black holes. The Reissner–Nordström (RN) black hole, a solution to the Einstein-Maxwell equations for a charged black hole, presents a unique challenge to these conjectures. If the charge $Q$ of the black hole exceeds its mass $M$ ($Q>M$), the RN solution would contravene the WCCC by revealing a naked singularity to distant observers. Conversely, when the RN black hole approaches an extremal state ($Q=M$), it suggests the existence of decay products with a charge-to-mass ratio exceeding unity. While this satisfies the WGC, it simultaneously breaches the WCCC, as these decay products cannot form black holes but are instead predicted to be elementary particles. This apparent contradiction between the WGC and the WCCC in the context of RN black holes is one of the intriguing puzzles at the intersection of quantum theory and gravitational physics[107, 108, 109].
The WGC and the WCCC are two distinct ideas in theoretical physics, each addressing different aspects of fundamental physics[110, 111, 112]. While both conjectures involve the term ”weak” and pertain to fundamental aspects of gravity and spacetime, they are distinct in their focus and implications. The WGC primarily concerns the relative strengths of forces in a quantum gravity framework, while the WCCC deals with the behavior of singularities and event horizons in gravitational collapse scenarios within classical general relativity. Combining these concepts, one might speculate about potential connections between the WGC and cosmic censorship, such as whether the WGC could have implications for cosmic censorship scenarios. In[113] researchers explore how Reissner-Nordström (R-N) black holes in perfect fluid dark matter (PFDM) might reconcile the weak gravity conjecture (WGC) and the cosmic censorship conjecture (WCCC). Normally, a charged black hole cannot violate the WCCC, but recent proposals suggest a connection between the WGC and WCCC. The study shows that with certain constraints, both conjectures can coexist. Without PFDM, an R-N black hole can lead to a naked singularity, violating the WCCC. However, with PFDM, the black hole always has event horizons, covering the singularity and satisfying both conjectures. A critical value of PFDM makes the black hole extreme, still fulfilling both WGC and WCCC. Thus, PFDM allows compatibility between the WGC and WCCC in R-N black holes. In this work, we intend to consider another approach to investigate the weak gravity conjecture and the cosmic censorship conjecture with regard to the thermodynamics of black holes in different structures such as EPST, RPST, and CFT. We aim to challenge these conjectures and the differences in these spaces and examine specific parameters such as AdS radii and central charge etc. Our investigations have revealed that under specific conditions, such as the incorporation of PFDM, these conjectures can indeed coexist harmoniously. This concord is vital as it ensures that our theoretical models remain consistent and do not yield contradictory predictions about the thermodynamic behavior of black holes. By demonstrating compatibility under these conditions, we contribute to a more coherent understanding of the fundamental principles governing black hole physics.